OpenStudy (anonymous):

Simplify 4√6/√30 by rationalizing the denominator

OpenStudy (anonymous):

multiply above and below by rooy30

OpenStudy (anonymous):

\[\LARGE \frac{4\sqrt6}{\sqrt{30}}\cdot \frac{\sqrt{30}}{\sqrt{30}}=...\] can you do it now ? :)

OpenStudy (anonymous):

4sqrt5/5 should be final answer..

OpenStudy (anonymous):

how do I get that?

OpenStudy (anonymous):

yeah... I'd also like to know ! :P

OpenStudy (anonymous):

@mrasksalotofquestions

OpenStudy (anonymous):

... I'll do it \[\LARGE \frac{4\sqrt6 \cdot \sqrt{30}}{(\sqrt{30})^2}=\frac{4\cdot \sqrt{6\cdot 30}}{30}\] \[\LARGE \frac{4\sqrt{180 }}{30}=\frac{4\cdot \sqrt{36 \cdot 5}}{30}=\] \[\LARGE \frac{4\sqrt{6^2}\cdot \sqrt5}{30}=\frac{4\cdot 6\cdot \sqrt5 }{30}=\] \[\LARGE \frac{24\cdot \sqrt5}{30}=\frac{4\cdot \sqrt5 }{5}\]

OpenStudy (anonymous):

thank you so much